# Low dimensional representation of multi-patient flow cytometry datasets using optimal transport for measurable residual disease detection in leukemia

### v2

@@ 2.5.1 Statistical analysis on the K-means weight vectors
As the measures $(\nu^{i})_{i=1}^{N}$ share a common support, it is natural to compare them in terms of their weight vectors $a^{i}=(\mu^{i}(V_{1}),\ldots,\mu^{i}(V_{K}))_{i=1}^{N}$, belonging to the probability simplex $\Sigma_{K}$. We can then apply the classical tools of statistical analysis of compositional data belonging to the probability simplex [3].
@@ 2.5.2 Statistical analysis with linearized OT
The new methodology that we propose to represent and cluster FCM datasets consists in:
@@
    Computing a Wasserstein barycenter $\nu_{\text{bar}}$ (4) of the measures $(\nu^{i})_{i=1}^{N}$, with a support of size $K$.
@@ 2.6 Performance evaluation
For some cytometry datasets of the DATAML-Bordeaux study, ground truth labels are available with the measures (MRD-BioM), giving a percentage linked to quantities expressing the presence of blasts. MRD-BioM is then defined as being positive when its value is higher than $0.1\%$ and negative otherwise. We will then refer to positive and negative follow-ups respectively. On the other hand, the FlowSOM method also provides a measure of MRD, MRD-FlowS, which is related to the percentage of blasts in the bone marrow. In cytometry, the threshold for positivity is also defined as $0.1\%$.
@@ 3.1.1 PCA for visualization
    Log-ratio transform of the compositional data $(a^{i}_{k})_{i=1}^{N}$ (K-Means + comp)
@@
We observe that the PCA visualization of the linearized data following the preliminary K-means step is highly dependant on the number K of clusters. Different experiments show a nuanced relationship between cluster quantity and discriminative power. As the number of clusters increases, the datasets’ distinguishable features gradually fade away. In [40], the authors opted for K = 64 nodes in the FlowSOM grid, a decision that echoes our own findings as a reasonable choice. For the KME, two parameters are required: the bandwidth $\sigma$ for RBF kernel and the dimension $s$ of the Fourier features. In order to fairly compare results, we choose $s=K\times d=448$ as the linearization sends measures to vectors in $\mathbb{R}^{64\times 7}=\mathbb{R}^{448}$. Regarding the bandwidth of the kernel, we set $\sigma=5$, which allows to best discriminate the patients using PCA.
@@ 3.1.2 Silhouette scores
In Table 1, we report silhouette scores to evaluate how well each of the four embedding discriminates between clusters. We note that silhouette scores should not be compared for different dimensions: as distances become similar in higher dimensions, the score naturally worsen in those cases. Kernel mean embedding performs poorly and K-means followed by compositional analysis gives bad results when $K>16$. As expected from the PCA visualisations, the silhouette scores for OT linearization on FlowSOM and K-means are highly similar.
Table 1: Silhouette scores for the different embeddings on HIPC. For each method, we compute the score either directly on raw data or on the data projected onto the first two principal components of the PCA. The best score per row is indicated in bold.
@@ 3.1.3 Execution time
Computations were performed using HPC resources from the MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux. In Figure 2, we compare the computational times of each embedding. K-means takes the upper hand, being two to three times faster than FlowSOM. KME is however significantly slower. The linearization of optimal transport takes less than 2 seconds no matter the dimension. The most time-consuming step in our process is computing the reference measure $\nu_{\text{ref}}$, which we have defined as the barycenter of all $N$ measures $(\nu_{i})_{i=1}^{N}$. Performing a Wasserstein barycenter is indeed costly, especially as the number K of clusters increases. However, for our dimensionality reduction purpose, K is kept small. In any case, one could set the reference measure as the uniform measure supported on the K-means centers $\nu_{\text{ref}}=\frac{1}{K}\sum_{k=1}^{K}\delta_{x_{k}}$. This procedure is significantly faster, but leads to a slight drop in the silhouette score.
@@ 3.2.2 PCA
In Figure 4, we compare two-dimensional representations of the dataset obtained by a PCA computed from the KME, the log-ratio transform of compositional data of K-means and the linearization of OT on K-means clustering. As for the HIPC datasets, we compare the PCA visualization on the linearized data for different number of clusters for the K-means step. This leads us to choose $K=64$. The OT linearization allows to visually discriminate diagnosis, normal bone marrow and follow-ups. It is however harder to distinguish positive from negative follow-ups.
In Figure 5, we visualize the amount of MRD in the follow-ups for the K-Means + LinW2 PCA method, where the size of the follow-up points represents the MRD-BioM value for the patient at the time that the sample was collected. We observe that high MRD-BioM follow-ups are pulled towards the set of diagnosis. Regarding the negative follow-ups, these are evidently more directed towards NBM and clearly separated from the set of diagnosis. We also visualize the MRD-FlowS predictions with the FlowSOM method in Figure 6, where we observe the same effect: high MRD-FlowS points lay in the diagnosis set whereas lower MRD-FlowS are closer to normal bone marrow points. Our low-dimensional representation of multi-patient FCM datasets is therefore consistent with the values of MRD-FlowS.
Figure 4: Two-dimensional representations of the DATAML-Bordeaux datasets using PCA on different embeddings : (top) K-Means + LinW2, (middle) K-Means + comp, (bottom) KME. Each dot represents a FCM sample. The colors encore the nature of the data : (red) diagnostic, (purple) normal bone marrow, (yellow) positive follow-up, (green) negative follow-up, (grey) no information on the MRD-BioM.
@@ 3.3 Silhouette scores
We finally report in Table 2 silhouette scores to evaluate how the various embeddings of the DATAML-Bordeaux dataset displayed in Figure 4 allow to discriminate clusters. We consider the following clusters: diagnosis, normal bone marrow, positive MRD-BioM follow-ups, and negative MRD-BioM follow-ups. We leave out the follow-ups for which we do not have a value MRD-BioM. From the results reported in Table 2, it can be seen that the clustering obtained with OT linearization is still the best one.
Table 2: Silhouette scores for the different embeddings on DATAML-Bordeaux. For each method, we compute the score either directly on raw data or on the data projected onto the first two principal components of the PCA. The best score per row is indicated in bold.
@@
None -0.19 -0.21
16 -0.03 -0.07 0.03 0.1
32 -0.1 -0.06 0.04 0.19
64 -0.03 0.04 0.03 0.17
128 -0.01 0.15 0.03 0.19
256 -0.01 0.15 0.02 0.19
@@ 4 Discussion
In this study, we have introduced a novel approach, via mean measure quantization and linearized OT, for low-dimensional embedding in a linear space of multi-patient FCM datasets viewed as discrete probability measures with large dimensional supports. We assessed the benefits of OT linearization over compositional data analysis and kernel mean embedding by showing that our OT-based approach yields the best silhouette scores for patients clustering. The biological relevance of our OT-based embedding has also been assessed by relating the consistency of our low-dimensional representation of multi-patient FCM datasets to the level of estimated MRD in follow-up using either molecular biology or the FlowSOW method. Interestingly, patients with positive follow-ups and large MRD values tend to be clustered, in our low-dimensional representation, with data points representing FCM from diagnosis. These results suggest that using OT for statistical learning from multi-patient FCM datasets is an interesting research direction in order to help the experts with MRD assessment. Indeed, combined with the FlowSOM method, our approach could enhance the accuracy of the relapse prognosis which is critical for making informed treatment decisions for AML patients.
@@ 5 Acknowledgments
This work benefited from financial support from the French government managed by the National Agency for Research under the France 2030 program, with the reference ANR-23-PEIA-0004. We would like to thank the data management unit of Toulouse University Hospital and the CAPTOR (Cancer Pharmacology of Toulouse Oncopole and Region) project (ANR-11-PHUC-001) for his financial support enabling e-CRF.

### v3

@@ 2.5.1 Statistical Analysis on the K-means Weight Vectors
As the measures $(\nu^{i})_{i=1}^{N}$ share a common support, it is natural to compare them in terms of their weight vectors $a^{i}=(\mu^{i}(V_{x_{1}}),\ldots,\mu^{i}(V_{x_{K}}))_{i=1}^{N}$, belonging to the probability simplex $\Sigma_{K}$. We can then apply the classical tools of statistical analysis of compositional data belonging to the probability simplex [3].
@@ 2.5.2 Statistical Analysis with Linearized OT
The new methodology that we propose to represent and cluster FCM datasets consists of:
@@
    Computing a Wasserstein barycenter $\nu_{\text{bar}}$ 4 of the measures $(\nu^{i})_{i=1}^{N}$, with a support of size $K$.
@@ 2.6 Performance Evaluation
For some cytometry datasets of the DATAML-Bordeaux study, ground truth labels are available with the measures (MRD-BioM), giving a percentage linked to quantities expressing the presence of blasts. MRD-BioM is then defined as being positive when its value is higher than $0.02\%$ and negative otherwise. We will then refer to positive and negative follow-ups respectively. On the other hand, the FlowSOM method also provides a measure of MRD, MRD-FlowS, which is related to the percentage of blasts in the bone marrow. Recent studies [38, 47] demonstrate that the commonly used 0.1% threshold in flow cytometry may not optimally differentiate patients with low versus high relapse risk. These studies identify an optimal prognostic threshold in the range of 0.01% to undetectable levels. Based on these findings, we define MRD positivity as the detection of at least 20 LAIP events within a 500,000-cell sample, consistent with established statistical confidence thresholds for flow cytometry.
@@ 3.1.1 PCA for Visualization
    Log-ratio transform of the compositional data $(a^{i})_{i=1}^{N}$ (K-Means + comp)
@@
When selecting the number of clusters $K$, one aims at compromising between performance and computational efficiency. In Figure 1, we compare the silhouette score (representing performance) and the execution time of our method on the HIPC dataset when $K$ varies from 10 to 250. The silhouette score is computed in the three-dimensional space obtained after a three-component PCA applied on the LOT embedded data. We chose to keep three components with an elbow criterion. We observe that for $K\geq 50$, the silhouette score stabilizes around 0.66 whereas the time of execution grows rapidly. Moreover, the authors of the FlowSOM method [43] chose $K=64$ for the number of clusters, which in our analysis corresponds to a good trade-off between performance and computational time. In the same Figure 1, we also compare different choices for the reference measure in the LOT embedding (10). The first natural choice is the Wasserstein barycenter $\nu_{\mathrm{bar}}$ (4) of the $\nu^{i}$’s. We also consider as a reference measure the uniform distribution supported on the the K-means centers $\nu_{\mathrm{unif}}=(1/K)\sum_{k=1}^{K}\delta_{x_{k}}$ and a measure $\nu_{\mathrm{rand}}=\nu^{i}$ randomly chosen among the $\nu^{i}$’s. Another possibility introduced in [25] is to choose multiple references for multiple embeddings. We also tested this option by randomly choosing two indexes $i,j$ and performing LinW2 with respect to the product space $L^{2}(\nu^{i})\times L^{2}(\nu^{j})$. We observe that choosing the Wasserstein barycenter as the reference measure performs better in general. This could be computationally expensive but remains reasonable when $K=64$.
Figure 1: HIPC dataset : Silhouette score and execution time of our method as a function of the number of clusters. The red dotted line represents the execution time for the reference measure chosen as either uniform, random among the $\nu^{i}$’s or the concatenation of two measures $(\nu^{i},\nu^{j})$.
For the KME, two parameters are required: the bandwidth $\sigma$ for RBF (Radial Basis Function) kernel and the dimension $s$ of the Fourier features. In order to fairly compare results, we choose $s=K\times d=448$ as the linearization sends measures to vectors in $\mathbb{R}^{64\times 7}=\mathbb{R}^{448}$. Regarding the bandwidth of the kernel, we set $\sigma=5$, which allows to best discriminate the patients using PCA.
@@ 3.1.2 Silhouette Scores
In Table 1, we report silhouette scores to evaluate how well each of the four embedding discriminates between clusters. We note that silhouette scores should not be compared for different dimensions: as distances become similar in higher dimensions, the score naturally worsen in those cases. Kernel mean embedding performs poorly and K-means followed by compositional analysis gives bad results when $K>16$. As expected from the PCA visualizations, the silhouette scores for OT linearization on FlowSOM and K-means are highly similar.
Table 1: Silhouette scores for the different embeddings on HIPC. For each method, we compute the score either directly on raw data or on the data projected onto the first three (chosen with the elbow criterion) principal components of the PCA. The best score per row is indicated in bold.
@@ 3.1.3 Execution Time
Computations were performed using HPC resources from the MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux. In Figure 3, we compare the computational times of each embedding. K-means takes the upper hand, being two to three times faster than FlowSOM. KME is however significantly slower. The linearization of optimal transport takes less than 2 seconds no matter the dimension. The most time-consuming step in our process is computing the reference measure $\nu_{\text{ref}}$, which we have defined as the barycenter of all $N$ measures $(\nu^{i})_{i=1}^{N}$. Performing a Wasserstein barycenter is indeed costly, especially as the number K of clusters increases. However, for our dimensionality reduction purpose, K is kept small. In any case, one could set the reference measure as the uniform measure supported on the K-means centers $\nu_{\text{ref}}=\frac{1}{K}\sum_{k=1}^{K}\delta_{x_{k}}$. This procedure is significantly faster, but leads to a slight drop in the silhouette score.
@@ 3.2.2 PCA
In Figure 5, we compare two-dimensional representations of the dataset obtained by a PCA computed from the KME, the log-ratio transform of compositional data of K-means and the linearization of OT on K-means clustering. The OT linearization allows us to visually discriminate between diagnosis, normal bone marrow and follow-ups. It is however harder to distinguish positive from negative follow-ups.
In Figure 6, we visualize the amount of MRD in the follow-ups for the K-Means + LinW2 PCA method, where the size of the follow-up points represents the MRD-BioM value for the patient at the time that the sample was collected. We observe that high MRD-BioM follow-ups are pulled towards the set of diagnoses. Regarding the negative follow-ups, these are evidently more directed towards NBM and clearly separated from the set of diagnoses. We also visualize the MRD-FlowS predictions with the FlowSOM method in Figure 7, where we observe the same effect: high MRD-FlowS points lie in the diagnosis set whereas lower MRD-FlowS are closer to normal bone marrow points. Our low-dimensional representation of multi-patient FCM datasets is therefore consistent with the values of MRD-FlowS.
Figure 5: Two-dimensional representations of the DATAML-Bordeaux datasets using PCA on different embeddings : (top) K-Means + LinW2, (middle) K-Means + comp, (bottom) KME. Each dot represents a FCM sample. The colors encore the nature of the data : (red) diagnostic, (purple) normal bone marrow, (yellow) positive follow-up, (green) negative follow-up, (gray) no information on the MRD-BioM.
@@ 3.3 Silhouette Scores
We finally report in Table 2 silhouette scores to evaluate how the various embeddings of the DATAML-Bordeaux dataset displayed in Figure 5 allow us to discriminate clusters. We consider the following clusters: diagnosis, normal bone marrow, positive MRD-BioM follow-ups, and negative MRD-BioM follow-ups. We leave out the follow-ups for which we do not have a value for MRD-BioM. From the results reported in Table 2, it can be seen that the clustering obtained with OT linearization is still the best one.
Table 2: Silhouette scores for the different embeddings on DATAML-Bordeaux. For each method, we compute the score either directly on raw data or on the data projected onto the first three (chosen by the elbow criterion) principal components of the PCA. The best score per row is indicated in bold.
@@
None -0.19 -0.22
16 -0.015 -0.2 0.04 0.04
32 -0.09 -0.04 0.07 0.09
64 -0.004 -0.1 0.06 0.1
128 -0.03 0 0.06 0.08
256 -0.02 0.02 0.04 0.08
3.4 Supervised Classification
One can leverage our dimension reduction method for classification purposes. After applying a three-component PCA on the different embeddings, we consider two classes : positive follow-ups (Fu+) and negative follow-ups (Fu-). For each embedding, we apply a three-component PCA on the embedded data, train a logistic regression model and predict either a positive or negative follow-up with a leave-one-out cross validation technique. We obtain the confusion matrices shown in Figure 8. In order to compare the different methods, we also provide in Figure 8 the balanced accuracy scores, which allow evaluating classifier performances when dealing with imbalanced data, that is when one of the classes is more present than the other. We observe that the LOT method outperforms the others in terms of this score.
Figure 8: Confusion matrices and balanced accuracy scores for the logistic regression classifier on the (left) K-means+LinW2+PCA, (middle) K-means+Comp+PCA and (right) KME+PCA models. The balanced accuracy score is defined as the arithmetic mean between specificity and sensitivity : $\frac{1}{2}\bigl{(}\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\bigr{)}.$
@@ 4 Discussion
In this study, we have introduced a novel approach, via mean measure quantization and linearized OT, for low-dimensional embedding in a linear space of multi-patient FCM datasets viewed as discrete probability measures with large dimensional supports. We assessed the benefits of OT linearization over compositional data analysis and kernel mean embedding by showing that our OT-based approach yields the best silhouette scores for patient clustering. The biological relevance of our OT-based embedding has also been assessed by relating the consistency of our low-dimensional representation of multi-patient FCM datasets to the level of estimated MRD in follow-up using either molecular biology or the FlowSOW method. Interestingly, patients with positive follow-ups and large MRD values tend to be clustered, in our low-dimensional representation, with data points representing FCM from diagnosis. These results suggest that using OT for statistical learning from multi-patient FCM datasets is an interesting research direction in order to help the experts with MRD assessment. Combined with the FlowSOM method, our approach could enhance the accuracy of relapse prognosis which is critical for making informed treatment decisions for AML patients. Indeed, if biologists have predicted a positive MRD with their method and the corresponding sample is close to the diagnosis in the two-dimensional embedding space, they can confirm their prediction. On the other hand, if they have predicted negative MRD and the sample is among diagnosis, they might want to take another look at their analysis.
@@ 5 Acknowledgments
This work benefited from financial support from the French government managed by the National Agency for Research under the France 2030 program, with the reference ANR-23-PEIA-0004. We would like to thank the data management unit of Toulouse University Hospital and the CAPTOR (Cancer Pharmacology of Toulouse Oncopole and Region) project (ANR-11-PHUC-001) for its financial support enabling e-CRF.
Conflict of interest. The authors have no conflict of interest to declare.
Author contributions. Erell Gachon: conceptualization, methodology, writing - original draft, investigation, writing - review and editing. Jérémie Bigot: conceptualization, methodology, writing - original draft, investigation, writing - review and editing, supervision, funding acquisition. Elsa Cazelles: conceptualization, methodology, writing - original draft, investigation, writing - review and editing, supervision, funding acquisition. Audrey Bidet : resources, data curation, validation, writing - review and editing. Jean-Philippe Vial: resources, data curation, validation, writing - review and editing. Pierre-Yves Dumas: resources, data curation, validation, writing - review and editing. Aguirre Mimoun: resources, data curation, validation, writing - review and editing.